# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

### 07 Dec 2023 v2

@@ 1 Introduction
RDMA technology has found extensive deployment in modern data centers [19, 10], offering low latency and high throughput benefits that are particularly advantageous for LLM training. RDMA technology primarily comprises two categories: InfiniBand and RoCE (RDMA over Converged Ethernet) [7, 12, 10]. Both of these technologies facilitate direct access to remote computer memory, bypassing the need for CPU intervention, thus leading to more efficient data transfer and communication. InfiniBand represents a dedicated networking technology, while RoCE is a protocol designed to implement RDMA within the Ethernet framework. Both InfiniBand and RoCE play an integral role in the construction of high-performance computing and data center networks [16, 12, 10]. Despite their widespread applications, the two networking technologies, InfiniBand and RoCE, are inherently incompatible with each other. This incompatibility poses challenges, especially in the context of training LLMs within clusters that incorporate various RDMA NICs. In such scenarios, the adoption of pure Ethernet for communication often emerges as the only viable option.
@@
In this paper, we present Holmes, a high-performance framework optimized for training LLMs in the heterogeneous RDMA NIC environment. An overview of the framework is illustrated in Figure 1. Holmes optimally leverages the diverse capabilities of NICs within cluster configurations. A key contribution lies in our intelligent scheduling methodology, strategically allocating computational tasklets to GPUs based on their associated NIC types. This approach aligns computational tasks with the strengths of each network technology, including InfiniBand, RoCE, and Ethernet. As a result, Holmes significantly mitigates overall training costs.
@@
The proposed Holmes framework demonstrates exceptional scalability through pipeline model parallelism, seamlessly extending to multiple GPU clusters. This scalability is maintained even in scenarios where high-speed interconnects (InfiniBand or RoCE) are unavailable between nodes from distinct clusters. Therefore, Holmes can be applied to train LLMs across multiple GPU clusters at different locations. This maximizes the utilization of existing GPU clusters for LLM training, eliminating the need to reconstruct larger GPU clusters to accommodate numerous GPU devices and RDMA NICs specifically for training purposes. To validate Holmes’s scalability, we conducted an extensive range of experiments covering a diverse spectrum of scenarios involving heterogeneous NICs. Remarkably, our training framework consistently achieved performance levels close to those attainable in environments with homogeneous RDMA-capable networks, such as pure InfiniBand or RoCE. Furthermore, Holmes notably surpassed training efficiency even in the pure Ethernet environment.
In general, Holmes provides a highly optimized framework for training LLMs in the heterogeneous RDMA NIC environment. Our optimizations of schedule and parallelism unlock the potential of heterogeneous systems for efficient and scalable LLM training. We firmly believe that Holmes signifies a significant step in democratizing LLMs, making efficient scaling attainable for a broader research community.
Our contributions are summarized as follows:
  • We propose Holmes, a novel LLM training framework designed to adapt to the heterogeneous NIC environment. Holmes is non-intrusive to existing infrastructures, harnessing pipeline model parallelism technologies. It facilitates LLM training across multiple GPU clusters without the need to reconstruct larger GPU clusters. This feature is crucial for conserving computational resources.
@@ 2.2 NIC Environments
  • Case 2: Heterogeneous Clusters with Absence of High-Speed Interconnects. These GPU devices are distributed among multiple clusters, each equipped with either the same or distinct types of RDMA NIC. Notably, there are no high-speed interconnects between these clusters. Communication via RDMA NICs is limited to GPU devices within the same cluster, whereas GPU devices located in different clusters communicate using Ethernet NICs.
@@ 3 Holmes Design
Figure 3: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices across the two clusters. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using NVLink.
@@ 3.1.1 Tensor Model Parallelism
${[\textbf{TP}]_{i,j}=rank_{(i-1)\cdot t+j}}\quad_{j\in\{1,2\ ,...,\ t\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ d\}}$ (1)
@@ 3.1.2 Pipeline Model Parallelism
Current LLM training frameworks exclusively support LLM training with low-speed Ethernet NICs within clusters that possess heterogeneous NICs. To address these limitations, we have introduced a solution to enable LLM training in the heterogeneous NIC environment. Capitalizing on the relatively low communication cost of pipeline parallelism, we establish pipeline parallelism groups among multiple clusters that lack high-speed interconnects. Consequently, these pipeline parallelism groups consist of GPU devices connected via heterogeneous NICs, and their sole means of communication is through Ethernet due to NIC incompatibility. Conversely, owing to the division of pipeline parallelism groups, the data parallelism groups can comprise GPU devices connected via homogeneous NICs, enabling communication via RDMA NICs. In terms of tensor parallelism, it is confined to the same node and communicates via NVLink interfaces.
@@
where $\alpha$ is the hyper-parameter. Similarly, the calculation of $N_{roce}$ follows the same method, ensuring $N_{ibz}+N_{roce}=N$. It can be easily extended to scenarios where the pipeline parallel size is more than 2. Our objective is to allocate a greater number of model layers to the GPU device connected to the faster NIC.
@@
${[\textbf{PP}]_{i,j}=rank_{i+(j-1)\cdot t\cdot d}}\quad_{j\in\{1,2\ ,...,\ p\}}^{i\in\{1,2\ ,...,\ t\ \cdot\ d\}}$ (3)
@@
3.2 Data Parallelism
  • Gradient Aggregation: At the end of each batch or epoch, GPUs share and combine their calculated gradient information using the Ring Allreduce algorithm [29]. This algorithm circulates gradient information from each GPU to others in a circular manner, accumulating gradients on each GPU. The process is repeated through multiple steps until all GPUs have received the global gradient information.
@@ 3.2 Data Parallelism
Automatic NIC Selection.
Figure 4: Code snippet of creating data parallel groups across clusters within the heterogeneous NIC environment.
Overlapped Distributed Optimizer.
@@
${[\textbf{DP}]_{i,j}=rank_{mod(i-1,t)+(floor(\frac{i-1}{t})\cdot d+j-1)\cdot t+1}}\ _{j\in\{1,2\ ,...,\ d\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ t\}}$ (4)
where $mod(i-1,t)$ denotes the modulo operation, which calculates the remainder when dividing two numbers, and $floor(*)$ is used to signify rounding down a real number.
@@ 4 Experiment
In this section, we seek to answer the following questions:
@@ 4.1 Experiment Setup
  • Hybird: There are multiple different clusters, each with nodes equipped with different high-speed NICs.
@@ 4.2 Performance Comparison
As discussed in 2.2, we compare two cases in the heterogeneous NIC environment. For Case 1 and Case 2, we evaluate Holmes across four types of NIC environments, including both homogeneous and heterogeneous high-speed interconnects, and varying numbers of nodes. The results are presented in Table 3, providing insights to address questions Q1 and Q2.
@@
Pipeline Partition.
Comparison with Other Training Frameworks.
@@ 5.2 Communication Optimization
Distributed training necessitates frequent parameter synchronization across nodes. In the initial stages of communication scheduling, the primary focus is on ensuring model consistency [33, 27, 13] and refining parameter synchronization architectures [18, 24, 11]. With the increasing scale of models, certain research endeavors explore coordinated optimization of both communication and parallel strategies [34, 40]. This involves implementing advanced strategies to enhance parallel efficiency while concurrently reducing communication demands.

### 11 Dec 2023 v3

@@ 1 Introduction
RDMA technology has found extensive deployment in modern data centers [19, 10], offering low latency and high throughput benefits that are particularly advantageous for LLM training. RDMA technology primarily comprises two categories: InfiniBand and RoCE (RDMA over Converged Ethernet) [7, 12, 10]. Both of these technologies facilitate direct access to remote computer memory, bypassing the need for CPU intervention, thus leading to more efficient data transfer and communication. InfiniBand represents a dedicated networking technology, while RoCE is a protocol designed to implement RDMA within the Ethernet framework. Both InfiniBand and RoCE play an integral role in the construction of high-performance computing and data center networks [16, 12, 10]. Despite their widespread applications, the two networking technologies, InfiniBand and RoCE, are inherently incompatible with each other. This incompatibility poses challenges, especially in the context of training LLMs within or across clusters that incorporate various RDMA NICs. In such scenarios, the adoption of pure Ethernet for communication often emerges as the only viable option.
@@
In this paper, we present Holmes, a high-performance framework optimized for training LLMs in the heterogeneous RDMA NIC environment. An overview of the framework is illustrated in Figure 1. Holmes optimally leverages the diverse capabilities of NICs within cluster configurations. A key contribution lies in our intelligent scheduling methodology, strategically allocating computational tasklets to GPUs based on their associated NIC types. This approach aligns computational tasklets with the strengths of each network technology, including InfiniBand, RoCE, and Ethernet. As a result, Holmes significantly mitigates overall training costs.
@@
The proposed Holmes framework demonstrates exceptional scalability through pipeline model parallelism, seamlessly extending to multiple GPU clusters. This scalability is maintained even in scenarios where high-speed interconnects (InfiniBand or RoCE) are unavailable between nodes from distinct clusters. Therefore, Holmes can be applied to train LLMs across multiple GPU clusters at different locations. This maximizes the utilization of existing GPU clusters for LLM training, eliminating the need to reconstruct larger GPU clusters to accommodate numerous GPU devices and RDMA NICs specifically for training purposes. To validate Holmes’s scalability, we conducted an extensive range of experiments covering a diverse spectrum of scenarios involving heterogeneous NICs. Remarkably, our training framework consistently achieved performance levels close to those attainable in environments with homogeneous RDMA-capable networks, such as pure InfiniBand or RoCE. Furthermore, Holmes notably surpassed training efficiency in the pure Ethernet environment.
In general, we firmly believe that Holmes signifies a notable advancement in democratizing LLMs, making efficient scaling attainable for a broader research community. Our contributions are summarized as follows:
  • We propose Holmes, a novel LLM training framework designed to adapt to the heterogeneous NIC environment. Holmes is non-intrusive to existing infrastructures, harnessing data and model parallelism technologies. It facilitates LLM training across multiple GPU clusters without the need to reconstruct larger GPU clusters. This feature is crucial for conserving computational resources.
@@ 2.2 NIC Environments
  • Case 2: Heterogeneous Clusters without High-Speed Interconnects. These GPU devices are distributed among multiple clusters, each equipped with either the same or distinct types of RDMA NIC. Notably, there are no high-speed interconnects between these clusters. Communication via RDMA NICs is limited to GPU devices within the same cluster, whereas GPU devices located in different clusters communicate using Ethernet NICs.
@@ 3 Holmes Design
Figure 3: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using NVLink.
@@ 3.1.1 Tensor Model Parallelism
${[\textbf{TP}]_{i,j}=\operatorname{rank}_{(i-1)\cdot t+j}}\quad_{j\in\{1,2\ ,...,\ t\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ d\}}$ (1)
@@ 3.1.2 Pipeline Model Parallelism
Current LLM training frameworks exclusively support LLM training with low-speed Ethernet NICs within clusters that possess heterogeneous NICs. To address these limitations, we have introduced a solution to enable LLM training in the heterogeneous NIC environment. Capitalizing on the relatively low communication cost of pipeline parallelism, we establish pipeline parallelism groups among multiple clusters that lack high-speed interconnects. Consequently, these pipeline parallelism groups consist of GPU devices connected via heterogeneous NICs, and their sole means of communication is through Ethernet due to NIC incompatibility. Conversely, owing to the division of pipeline parallelism groups, the data parallelism groups can comprise GPU devices connected via homogeneous NICs, enabling communication via RDMA NICs. In terms of tensor parallelism, it is confined to the same node and communicates via PCI-E or NVLink interfaces.
@@
where $\alpha$ is the hyper-parameter. Similarly, the calculation of $N_{roce}$ follows the same method, ensuring $N_{ib}+N_{roce}=N$. It can be easily extended to scenarios where the pipeline parallel size is more than 2. Our objective is to allocate a greater number of model layers to the GPU device connected to the faster NIC.
@@
${[\textbf{PP}]_{i,j}=\operatorname{rank}_{i+(j-1)\cdot t\cdot d}}\quad_{j\in\{1,2\ ,...,\ p\}}^{i\in\{1,2\ ,...,\ t\ \cdot\ d\}}$ (3)
@@
Figure 4: A code snippet for creating data parallel groups across clusters within the heterogeneous NIC environment.
@@
3.2 Data Parallelism
  • Gradient Aggregation: At the end of each batch or epoch, GPUs use the Ring Allreduce algorithm [29] to share and combine their calculated gradient information. This algorithm circulates gradient details in a circular manner among GPUs, accumulating gradients on each GPU until all receive the global gradient information.
Automatic NIC Selection.
Overlapped Distributed Optimizer.
@@ 3.2 Data Parallelism
${[\textbf{DP}]_{i,j}=\operatorname{rank}_{\operatorname{mod}(i-1,t)+(\operatorname{floor}(\frac{i-1}{t})\cdot d+j-1)\cdot t+1}}\ _{j\in\{1,2\ ,...,\ d\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ t\}}$ (4)
where $\operatorname{mod}(i-1,t)$ denotes the modulo operation, which calculates the remainder when dividing two numbers, and $\operatorname{floor}(*)$ is used to signify rounding down a real number.
@@ 4 Experiment
In this section, we seek to answer the following questions:
@@ 4.1 Experiment Setup
  • Hybird: There are multiple different clusters, each with nodes equipped with different high-speed NICs. The connection between clusters are restricted to Ethernet.
@@ 4.2 Performance Comparison
As discussed in Section 2.2, we compare two cases in the heterogeneous NIC environment. For Case 1 and Case 2, we evaluate Holmes across four types of NIC environments, including both homogeneous and heterogeneous high-speed interconnects, and varying numbers of nodes. The results are presented in Table 3, providing insights to address questions Q1 and Q2.
@@
Pipeline Partition.
Comparison with Other Training Frameworks.
@@ 5.2 Communication Optimization
Distributed training necessitates frequent parameter synchronization. In the initial stages of communication scheduling, the primary focus is on ensuring model consistency [33, 27, 13] and refining parameter synchronization architectures [18, 24, 11]. With the increasing scale of models, certain research endeavors explore coordinated optimization of both communication and parallel strategies [34, 40]. This involves implementing advanced strategies to enhance parallel efficiency while concurrently reducing communication demands.